for a gas Cv= 4.96 cal/mole-K when 2 mole gas is heated from 340 K to 342K increase in internal energy is:
(1)9.92 Cal     (2) 13.90 Cal      (3) 19.84 Cal       (4) 27.80 Cal  

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Internal energy change  = n * Cv * \DeltaT
 = 2 * 4.96 * 2 = 19.84 Cal
 
Regards
Arun (askIITians forum expert)